Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS September/October 2010  Not logged in ELOG logo
Message ID: 96     Entry time: Wed Oct 6 23:48:49 2010
Author: Markus Friedl 
Project: BELLE 
Subject: run 
Run Number: run052 
Events: 165789 
StartTime: 06.10.2010 23:47:12 
EndTime: 07.10.2010 08:43:42 

Belle stack without Origami2
HV=80V I=217uA
telescope setting 103 / 30 / 277.5
telescope run 20184

temperature of chiller set to 15°C

beam stopped at 7.10. 08:00 sharp, no beam until 8.10 - 8:00 according to SPS Page 1. We keep DAQ and telescope running and cross our fingers for at least some spills of particles.

 

Irradiated micron module 2, n-side: excessive hit count on channels at the edges of the APV chips - no such thing on the p-side

ELOG V3.1.4-966e3dd